# CodeCrash: Stress Testing LLM Reasoning under Structural and Semantic Perturbations

### v2

@@ 4.1 Experiment Settings
Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95), with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs. For o3-mini, we raise the limit to 20,000 tokens to avoid verbose reasoning. Due to resource constraints, we generate three candidates for direct inference and one for CoT inference per query; we provide a result stability analysis under $N=3$ in Appendix B. Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT inference (details provided in Appendix C). Furthermore, we adopt Pass@1 accuracy Chen et al. (2021) as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability. All perturbed results are reported as either absolute ($\Delta_{\text{\;{pp}}}=\text{Perturbed}-\text{VAN}$) or relative ($\Delta_{\%}=\frac{\text{Perturbed}-\text{VAN}}{\text{VAN}}\times 100\%$) differences from the corresponding vanilla baseline. Additionally, we use percent signs (%) to indicate relative differences for clarity in all tables.
4.2 Impact on Output Prediction
We first evaluate LLM robustness in output prediction tasks. Table 1 summarizes the absolute Pass@1 degradation for all models under direct inference across both datasets. Our findings in this subsection include:
Key Findings: LLMs demonstrate an inherent and multi-layered reliance on natural language information, including identifiers, contextual messages, and reasoning hints, which ultimately undermines their reliability in code reasoning and comprehension. LLMs exhibit limited robustness to structurally disorganized code, struggling to differentiate irrelevant code from essential logic. Aggregating multiple structural perturbations exacerbates this vulnerability, significantly impairing their reasoning reliability. Larger scales, newer versions, and domain-specific fine-tuning LLMs substantially enhance robustness, highlighting improved code understanding and reasoning capabilities.
Table 1: Absolute Pass@1 degradation of LLMs under all perturbation types for output prediction tasks on Crux and LCB using direct inference ($N=3$). VAN shows the Pass@1 accuracy (in %), others indicate absolute drops from VAN (in  pp). Darker red highlights represent more severe degradation. See [To Relative Version] for the relative version.
Model Series Model Name VAN REN RTF GBC ALL MDC MPS MHC
Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB Crux LCB
GPT-4 Omni GPT-4o-Mini $57.29$ $53.24$ $-0.62$ $-9.67$ $-1.79$ $-0.97$ $-10.92$ $-5.50$ $-15.96$ $-16.98$ $-18.67$ $-21.50$ $-13.29$ $-14.89$ $-6.62$ $-23.59$
GPT-4o $71.33$ $64.51$ $-2.71$ $-9.60$ $-3.38$ $-4.59$ $-9.21$ $-3.20$ $-10.71$ $-18.30$ $-9.96$ $-18.79$ $-12.29$ $-15.66$ $-4.54$ $-17.19$
Claude Series Claude-3.5-Haiku-20241022 $57.42$ $58.25$ $-3.96$ $-9.32$ $-1.79$ $+0.35$ $-7.29$ $-4.52$ $-13.79$ $-21.50$ $-7.92$ $-5.85$ $-6.46$ $-4.66$ $-15.79$ $-31.04$
Claude-3.5-Sonnet-20241022 $71.46$ $73.83$ $-1.67$ $-16.77$ $-0.92$ $+1.18$ $-6.88$ $-4.94$ $-10.54$ $-25.19$ $-5.79$ $-7.10$ $-6.17$ $-7.86$ $-10.29$ $-32.08$
Gemini Series Gemini-1.5-Flash $56.21$ $44.26$ $-3.54$ $-0.42$ $-1.17$ $-1.04$ $-4.58$ $-0.35$ $-10.33$ $-5.64$ $-8.29$ $-14.27$ $-16.00$ $-17.88$ $-4.12$ $-9.81$
Gemini-2.0-Flash $65.04$ $65.97$ $-2.33$ $-18.09$ $-2.08$ $-0.77$ $-15.50$ $-8.77$ $-21.63$ $-27.77$ $-17.42$ $-34.24$ $-21.54$ $-33.19$ $-5.38$ $-12.87$
Gemini-1.5-Pro-002 $67.37$ $56.16$ $-2.92$ $-3.41$ $-2.50$ $-4.73$ $-5.62$ $-3.83$ $-13.42$ $-13.01$ $-15.50$ $-18.51$ $-14.62$ $-18.44$ $-6.75$ $-18.23$
DeepSeek DeepSeek-V3 $67.87$ $67.78$ $-2.67$ $-16.35$ $-1.33$ $-3.76$ $-4.62$ $-5.98$ $-8.75$ $-24.15$ $-11.25$ $-28.04$ $-6.87$ $-23.17$ $-7.29$ $-20.11$
LLaMA Series LLaMA-3.1-8B-Instruct $35.96$ $34.66$ $+0.08$ $-4.31$ $-0.33$ $-1.67$ $-3.13$ $-4.73$ $-8.54$ $-6.82$ $-8.08$ $-9.88$ $-6.42$ $-8.00$ $-6.00$ $-13.57$
LLaMA-3.1-70B-Instruct $56.08$ $44.89$ $-2.38$ $-1.39$ $-0.54$ $-1.74$ $-5.38$ $-2.44$ $-11.00$ $-7.72$ $-10.79$ $-12.53$ $-14.96$ $-17.40$ $-4.71$ $-14.47$
LLaMA-3.1-405B-Instruct $63.54$ $50.66$ $-5.37$ $-3.90$ $-0.62$ $-4.45$ $-5.79$ $-2.85$ $-12.50$ $-8.56$ $-4.17$ $-7.45$ $-7.17$ $-10.09$ $-4.37$ $-13.01$
LLaMA-3.3-70B-Instruct $59.92$ $48.50$ $-2.21$ $-1.39$ $-1.25$ $-3.41$ $-5.50$ $-5.71$ $-10.37$ $-9.74$ $-7.46$ $-9.60$ $-8.08$ $-12.32$ $-4.08$ $-10.16$
Qwen Series Qwen2.5-7B-Instruct $43.29$ $41.41$ $-3.04$ $-7.24$ $-1.29$ $-3.20$ $-12.96$ $-3.76$ $-16.42$ $-12.80$ $-25.13$ $-15.87$ $-19.58$ $-8.21$ $-11.67$ $-23.03$
Qwen2.5-14B-Instruct $47.79$ $49.48$ $-1.83$ $-9.88$ $-0.67$ $-1.39$ $-11.21$ $-3.41$ $-19.04$ $-14.82$ $-16.67$ $-20.46$ $-15.54$ $-16.77$ $-4.46$ $-11.00$
Qwen2.5-32B-Instruct $60.04$ $59.64$ $-1.71$ $-10.79$ $-2.58$ $-0.49$ $-5.79$ $-1.39$ $-11.88$ $-20.81$ $-10.83$ $-19.49$ $-14.21$ $-14.06$ $-7.87$ $-18.37$
Qwen2.5-72B-Instruct $60.08$ $54.91$ $-1.04$ $-8.49$ $-1.87$ $+0.42$ $-10.71$ $-2.02$ $-14.00$ $-13.85$ $-10.21$ $-6.96$ $-14.46$ $-14.41$ $-9.63$ $-22.27$
Qwen2.5-Coder-32B-Instruct $66.96$ $56.65$ $-2.50$ $-8.42$ $-0.96$ $-2.02$ $-10.75$ $-2.78$ $-14.96$ $-15.80$ $-15.54$ $-20.74$ $-11.08$ $-17.19$ $-7.00$ $-12.32$
Average $-2.38$ $-8.20$ $-1.47$ $-1.90$ $-7.99$ $-3.89$ $-13.17$ $-15.50$ $-11.98$ $-15.96$ $-12.28$ $-14.95$ $-7.09$ $-17.83$
$\mathbf{-4.56}$ $\mathbf{-1.63}$ $\mathbf{-6.46}$ $\mathbf{-14.04}$ $\mathbf{-13.47}$ $\mathbf{-13.28}$ $\mathbf{-11.11}$
4.2.1 Over-reliance on Natural Language Semantics
We categorize NL dependence into three hierarchical levels: surface-level, contextual-level, and reasoning-level dependence.
Surface-level NL Dependence (REN).
Under the REN perturbation, models suffer more significant performance degradation on LCB ($8.20$pp on average) compared to Crux ($2.38$pp) due to the presence of meaningful function names in the vanilla settings of LCB, which are absent in Crux. High-performing models like Claude-3.5-Sonnet and DeepSeek-V3 suffer severe degradations exceeding $\mathbf{16}$pp on LCB, highlighting a significant reliance on semantically meaningful identifiers.
Contextual-level NL Dependence (MDC and MPS).
@@ 4.2.1 Over-reliance on Natural Language Semantics
Under contextual-level textual perturbations (MDC and MPS), LLMs exhibit severe performance drops ($13.47$pp and $13.28$pp, respectively). Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^2] If MDC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa. ($16/17$ models), implying the model sensitivity to misleading comments and prints primarily depends on dataset characteristics and (2) perturbation-method consistency [^3] If MDC causes greater degradation on Crux than LCB, so does MPS, and vice versa. ($14/17$ models), indicating an inherent cognitive preference for NL messages from specific formats. These phenomena reveal that LLMs inherently depend on embedded NL context and exhibit format-specific cognitive biases.
Reasoning-level NL Dependence (MHC).
MHC perturbation also induces substantial degradation ($11.11$pp on average), notably exacerbated in complex datasets (LCB: $17.83$pp vs. Crux: $7.09$pp), which is a complexity-amplified effect. Advanced models like Claude-3.5-Haiku and Sonnet experience over $30$pp performance degradation, far exceeding their degradation under surface or contextual perturbations. These results illustrate that (i) LLMs readily accept plausible hint comments as authoritative reasoning shortcuts, and (ii) this over-trust intensifies when task complexity surpasses the model’s reasoning capacity, which we revisit under CoT prompting in §4.3.1.
4.2.2 Low Resistance to Disorganized Coding Environments
From RTF perturbation, LLMs can effectively reason over logically equivalent but syntactically complex expressions with minimal degradation ($1.63$pp). However, the highest performance drop emerges under the GBC compared to REN and RTF, with an average degradation of $6.46$pp across both datasets. Notably, GBC is the only perturbation where degradation is more severe on LCB ($3.89$pp vs. $7.99$pp on Crux) due to uneliminated meaningful identifiers. This trend reverses under the ALL perturbation, where all structural perturbations are combined, leading to a greater performance drop on LCB ($15.50$pp) than on Crux ($13.17$pp). Additionally, we observe a compounding effect of structural perturbations under ALL perturbation, resulting in the most significant overall degradation ($14.04$pp), which surpasses any individual code structural perturbation effects.
4.2.3 Model Robustness Across Different Scales
@@
We further analyze reasoning reliability across different model scales and versions, comparing relative Pass@1 degradation under representative perturbations in Table 2. Due to length constraints, the table, weighting scheme, and additional evaluation setups are provided in Appendix D.2. We observe a clear and consistent trend except for Gemini families. Larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models, such as LLaMA (3.1-8B $<$ 70B $<$ 405B) and Qwen (2.5-7B $<$ 14B $<$ 32B $<$ 72B). Additionally, newer model versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) consistently exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively.
Table 2: Relative Pass@1 degradation under direct inference, aggregated over Crux and LCB. Weighted average uses: 1/2 ALL, 1/8 MDC, 1/8 MPS, and 1/4 MHC. [To Absolute Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
GPT-4 Omni GPT-4o-Mini 55.77 $-29.37\%$ $-35.51\%$ $-24.99\%$ $-23.83\%$ $-28.20\%$
GPT-4o 68.78 $-20.01\%$ $-19.64\%$ $-19.87\%$ $-13.96\%$ $-18.44\%$
Claude Series Claude-3.5-Haiku-20241022 57.73 $-28.85\%$ $-12.38\%$ $-10.03\%$ $-37.16\%$ $-26.52\%$
Claude-3.5-Sonnet-20241022 72.35 $-22.01\%$ $-8.67\%$ $-9.39\%$ $-25.28\%$ $-19.58\%$
Gemini Series Gemini-1.5-Flash 51.73 $-16.27\%$ $-21.30\%$ $-32.94\%$ $-12.89\%$ $-18.14\%$
Gemini-2.0-Flash 65.39 $-36.56\%$ $-36.19\%$ $-39.56\%$ $-12.48\%$ $-30.87\%$
Gemini-1.5-Pro-002 63.17 $-21.13\%$ $-26.73\%$ $-25.88\%$ $-18.43\%$ $-21.75\%$
DeepSeek DeepSeek-V3 67.84 $-21.41\%$ $-25.86\%$ $-19.14\%$ $-17.83\%$ $-20.79\%$
LLaMA Series LLaMA-3.1-8B-Instruct 35.47 $-22.23\%$ $-24.74\%$ $-19.81\%$ $-25.10\%$ $-22.96\%$
LLaMA-3.1-70B-Instruct 51.89 $-18.71\%$ $-22.49\%$ $-31.20\%$ $-17.33\%$ $-20.40\%$
LLaMA-3.1-405B-Instruct 58.72 $-18.63\%$ $-9.61\%$ $-14.51\%$ $-13.93\%$ $-15.81\%$
LLaMA-3.3-70B-Instruct 55.64 $-18.35\%$ $-15.20\%$ $-17.95\%$ $-12.11\%$ $-16.35\%$
Qwen Series Qwen2.5-7B-Instruct 42.59 $-35.30\%$ $-50.65\%$ $-35.72\%$ $-37.69\%$ $-37.87\%$
Qwen2.5-14B-Instruct 48.42 $-36.14\%$ $-37.30\%$ $-33.04\%$ $-14.16\%$ $-30.40\%$
Qwen2.5-32B-Instruct 59.89 $-25.44\%$ $-23.52\%$ $-23.63\%$ $-19.74\%$ $-23.55\%$
Qwen2.5-72B-Instruct 58.14 $-24.02\%$ $-15.37\%$ $-24.88\%$ $-25.21\%$ $-23.34\%$
Qwen2.5-Coder-32B-Instruct 63.10 $-24.42\%$ $-28.23\%$ $-21.72\%$ $-14.68\%$ $-22.12\%$
4.3 Impact of Reasoning on Model Robustness
This section examines how the reasoning process, including CoT prompting for LLMs and internal reasoning in LRMs, affects model robustness against perturbations.
4.3.1 Effect of CoT Prompting
@@
Table 3: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction tasks in direct inference ($N=3$) and CoT inference ($N=1$). [To Absolute Version]
Model Series Model Name VAN ALL ($1/2$) MDC ($1/8$) MPS ($1/8$) MHC ($1/4$) Weighted Average
@@
GPT-4 Omni GPT-4o-Mini $55.77$ $81.47$ $-29.30\%$ $-13.15\%$ $-35.37\%$ $-10.84\%$ $-24.91\%$ $-10.94\%$ $-23.27\%$ $-2.69\%$ $-28.00\%$ $-9.52\%$
GPT-4o $68.78$ $91.79$ $-19.70\%$ $-4.94\%$ $-19.29\%$ $-5.03\%$ $-19.70\%$ $-5.62\%$ $-13.49\%$ $-1.36\%$ $-18.10\%$ $-4.22\%$
Claude Series Claude-3.5-Haiku-20241022 $57.73$ $72.95$ $-28.89\%$ $-21.33\%$ $-12.37\%$ $-10.61\%$ $-10.02\%$ $-8.68\%$ $-37.25\%$ $-14.26\%$ $-26.56\%$ $-14.30\%$
Claude-3.5-Sonnet-20241022 $72.35$ $86.00$ $-22.15\%$ $-7.55\%$ $-8.68\%$ $-5.27\%$ $-9.40\%$ $-6.36\%$ $-25.50\%$ $-7.91\%$ $-19.71\%$ $-6.86\%$
Gemini Series Gemini-1.5-Flash $51.73$ $75.22$ $-16.57\%$ $-18.30\%$ $-20.35\%$ $-21.73\%$ $-32.29\%$ $-42.20\%$ $-12.09\%$ $-2.08\%$ $-17.89\%$ $-20.39\%$
Gemini-2.0-Flash $65.39$ $89.13$ $-36.59\%$ $-6.23\%$ $-36.27\%$ $-6.32\%$ $-39.62\%$ $-14.30\%$ $-12.51\%$ $-2.02\%$ $-30.91\%$ $-7.01\%$
Gemini-1.5-Pro-002 $63.17$ $87.18$ $-21.00\%$ $-7.44\%$ $-26.32\%$ $-11.93\%$ $-25.41\%$ $-14.71\%$ $-17.49\%$ $-3.86\%$ $-21.34\%$ $-9.19\%$
DeepSeek DeepSeek-V3 $67.84$ $89.52$ $-21.40\%$ $-9.96\%$ $-25.85\%$ $-17.73\%$ $-19.13\%$ $-18.78\%$ $-17.83\%$ $-3.49\%$ $-20.78\%$ $-12.07\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.47$ $44.72$ $-22.26\%$ $-27.62\%$ $-24.69\%$ $-20.98\%$ $-19.76\%$ $-21.50\%$ $-24.91\%$ $-8.74\%$ $-22.92\%$ $-20.03\%$
LLaMA-3.1-70B-Instruct $51.89$ $69.43$ $-18.83\%$ $-17.91\%$ $-22.05\%$ $-23.20\%$ $-30.59\%$ $-30.86\%$ $-16.12\%$ $-6.76\%$ $-20.03\%$ $-19.22\%$
LLaMA-3.1-405B-Instruct $58.72$ $78.42$ $-18.77\%$ $-13.56\%$ $-9.19\%$ $-10.27\%$ $-14.07\%$ $-15.55\%$ $-12.96\%$ $-7.48\%$ $-15.53\%$ $-11.75\%$
LLaMA-3.3-70B-Instruct $55.64$ $76.86$ $-18.22\%$ $-10.89\%$ $-14.85\%$ $-7.73\%$ $-17.38\%$ $-11.09\%$ $-11.43\%$ $-4.78\%$ $-16.00\%$ $-8.70\%$
Qwen Series Qwen2.5-7B-Instruct $42.59$ $58.17$ $-35.37\%$ $-21.77\%$ $-50.86\%$ $-21.91\%$ $-35.99\%$ $-27.96\%$ $-37.39\%$ $-6.99\%$ $-37.89\%$ $-19.50\%$
Qwen2.5-14B-Instruct $48.42$ $70.68$ $-36.06\%$ $-23.12\%$ $-37.35\%$ $-35.29\%$ $-33.05\%$ $-43.36\%$ $-14.26\%$ $-10.84\%$ $-30.40\%$ $-27.34\%$
Qwen2.5-32B-Instruct $59.89$ $79.44$ $-25.41\%$ $-17.42\%$ $-23.50\%$ $-18.60\%$ $-23.63\%$ $-24.21\%$ $-19.71\%$ $-5.31\%$ $-23.53\%$ $-16.19\%$
Qwen2.5-72B-Instruct $58.14$ $82.80$ $-23.98\%$ $-12.65\%$ $-15.46\%$ $-6.80\%$ $-24.83\%$ $-16.24\%$ $-24.70\%$ $-6.61\%$ $-23.20\%$ $-10.63\%$
Qwen2.5-Coder-32B-Instruct $63.10$ $84.75$ $-24.20\%$ $-14.67\%$ $-27.72\%$ $-19.00\%$ $-21.19\%$ $-19.93\%$ $-14.25\%$ $-9.32\%$ $-21.78\%$ $-15.49\%$
Average $\mathbf{-24.63}\%$ $\mathbf{-14.62}\%$ $\mathbf{-24.13}\%$ $\mathbf{-14.90}\%$ $\mathbf{-23.59}\%$ $\mathbf{-19.55}\%$ $\mathbf{-19.72}\%$ $\mathbf{-6.15}\%$ $\mathbf{-23.21}\%$ $\mathbf{-13.67}\%$

### v3

@@ 3.1 Experiment Settings
Following LCB, we use nucleus sampling (temperature=0.2, top-p=0.95) with a maximum of 200 tokens for direct inference and 2,000 for CoT prompting, discarding excessively long outputs. Due to resource constraints, we generate three candidates for direct inference and one for CoT inference; we provide a result stability analysis under $N=3$ in Appendix C. Following Crux, we employ 2-shot prompting for direct inference and 1-shot for CoT step-by-step execution (details provided in Appendix D). Furthermore, we adopt Pass@1 accuracy as the primary evaluation metric and scale all scores in $[0,1]$ to percentages by multiplying by $100$ for readability. The experimental results are reported using relative differences from the corresponding vanilla baseline ($\Delta_{\%}=\frac{\text{Perturbed}-\text{{VAN}}}{\text{{VAN}}}\times 100\%$). We also report the absolute differences for completeness in the Appendix E.
3.2 Severe Degradation under Perturbations
We begin by evaluating models in the direct inference setting to measure their robustness against perturbations in §2.2 and summarize the results in Table 1. On average, models suffer a $\mathbf{23.2\%}$ performance drop, with the aggregated structural perturbation (PSC-ALL) causing the largest degradation ($24.6\%$). This indicates that LLMs exhibit limited code-tracing robustness, struggling to identify important code from essential logic and maintain reliability.
LLMs are sensitive to embedded NL cues.
@@ 3.2 Severe Degradation under Perturbations
As shown in Table 1, contextual-level perturbations (MCC and MPS) degrade performance by about $24\%$ on average, comparable to PSC-ALL, highlighting their substantial impact on model robustness. This also reveals that models are unable to consistently prioritize program logic over misleading NL cues, instead naively incorporating those cues into their reasoning. Notably, we identify two prominent consistency patterns: (1) dataset-level consistency [^1] If MCC degrades performance more on Crux than LCB, the same holds for MPS, and vice versa. ($15/17$ models), implying the impact intensity and sensitivity of both injection formats primarily depend on the dataset characteristics; and (2) perturbation-method consistency [^2] If MCC causes greater degradation on Crux than LCB, so does MPS, and vice versa. ($14/17$ models), indicating format-specific cognitive preference and biases. These phenomena suggest that LLMs are inherently impacted by embedded NL context, regardless of the injection format.
LLMs rationalize the hint to shortcut their reasoning.
Moving forward to the reasoning-level perturbation, MHC induces significant degradation with a complexity-amplified effect: models average drop $\mathbf{33.0\%}$ on LCB compared to $\mathbf{12.4\%}$ on Crux. Since LCB consists of more complex algorithmic problems with relatively simple output structures, models are more willing to adopt the hint as a reasoning shortcut. Notably, the performances of Claude-3.5-Haiku and Sonnet drop over $\mathbf{40\%}$ under MHC perturbation on LCB, which aligns with recent findings from Anthropic that models often accept a biased multiple-choice hint and rationalize it rather than critically resolving the contradiction Chen et al. (2025b). We extend this observation to the code reasoning domain and illustrate that: (i) LLMs readily rationalize the plausible hint as authoritative reasoning shortcuts, and (ii) this over-trust amplifies when task complexity surpasses the model’s reasoning capacity.
Scaling and versioning systematically improve robustness, with notable exceptions.
@@
Overall, robustness against perturbations improves consistently with model scale and newer releases. We observe a clear and consistent trend that larger closed-sourced models from GPT-4 Omni and Claude families demonstrate enhanced robustness compared to their smaller counterparts. Similar trends appear in open-source models, such as LLaMA (3.1-8B $\leq$ 70B << 405B) and Qwen (2.5-7B << 14B << 32B $\leq$ 72B). Additionally, newer versions (LLaMA-3.3-70B) and domain-specific fine-tuned variants (Qwen2.5-Coder-32B) exhibit improved robustness compared to their predecessors (LLaMA-3.1-70B) and larger counterparts (Qwen2.5-72B), respectively. Interestingly, the Gemini series deviates from this trend, underscoring that architectures and training strategies could affect robustness.
3.3 Limits of CoT in Resolving NL-Embedded Contradictions
@@
Table 2: Comparison of relative Pass@1 degradation aggregated over Crux and LCB for output prediction in different reasoning modes. See [Absolute Version] for the absolute degradation.
Model Series Model Name VAN PSC-ALL MCC MPS MHC Average
@@
GPT-4 Omni GPT-4o-Mini $55.8$ $81.5$ $29.4\%$ $13.1\%$ $35.5\%$ $10.8\%$ $25.0\%$ $11.0\%$ $23.8\%$ $2.7\%$ $28.4\%$ $9.4\%$
GPT-4o $68.8$ $91.8$ $20.0\%$ $4.9\%$ $19.6\%$ $5.0\%$ $19.9\%$ $5.6\%$ $0\% 1.4-1.4\%$ $18.4\%$ $4.2\%$
Claude Series Claude-3.5-Haiku-20241022 $57.7$ $72.9$ $28.9\%$ $21.2\%$ $12.4\%$ $6\% 10.0-1$$0\%$ $8.7\%$ $37.2\%$ $14.2\%$ $22.1\%$ $13.7\%$
Claude-3.5-Sonnet-20241022 $72.3$ $86.0$ $22.0\%$ $7.4\%$ $8.7\%$ $5.3\%$ $9.4\%$ $6.3\%$ $25.3\%$ $7.8\%$ $16.3\%$ $6.7\%$
Gemini Series Gemini-1.5-Flash $51.7$ $75.2$ $16.3\%$ $18.3\%$ $21.3\%$ $21.6\%$ $32.9\%$ $42.1\%$ $9\% 2.1-2.1\%$ $20.8\%$ $21.0\%$
Gemini-2.0-Flash $65.4$ $89.1$ $36.6\%$ $6.2\%$ $2\% 6.3-6.3\%$ $39.6\%$ $14.1\%$ $12.5\%$ $2.0\%$ $31.2\%$ $7.1\%$
Gemini-1.5-Pro-002 $63.2$ $87.2$ $21.1\%$ $7.4\%$ $26.7\%$ $11.9\%$ $25.9\%$ $14.6\%$ $18.4\%$ $3.9\%$ $23.0\%$ $9.4\%$
DeepSeek DeepSeek-V3 $67.8$ $89.5$ $21.4\%$ $9.9\%$ $25.9\%$ $17.6\%$ $19.1\%$ $\% 17.8-17.8\%$ $3.5\%$ $21.1\%$ $12.4\%$
LLaMA Series LLaMA-3.1-8B-Instruct $35.5$ $44.7$ $22.2\%$ $27.6\%$ $24.7\%$ $21.1\%$ $19.8\%$ $21.4\%$ $25.1\%$ $9.0\%$ $23.0\%$ $19.8\%$
LLaMA-3.1-70B-Instruct $51.9$ $69.4$ $18.7\%$ $18.0\%$ $22.5\%$ $23.3\%$ $2\% 31.1-3$$1\%$ $17.3\%$ $6.8\%$ $22.4\%$ $19.8\%$
LLaMA-3.1-405B-Instruct $58.7$ $78.4$ $18.6\%$ $13.6\%$ $9.6\%$ $10.3\%$ $14.5\%$ $15.6\%$ $13.9\%$ $7.5\%$ $4.2\% 11$$1.8\%$
LLaMA-3.3-70B-Instruct $55.6$ $76.9$ $18.4\%$ $10.8\%$ $15.2\%$ $7.7\%$ $7.9\% 11$$1.0\%$ $12.1\%$ $4.7\%$ $15.9\%$ $8.6\%$
Qwen Series Qwen2.5-7B-Instruct $42.6$ $58.2$ $35.3\%$ $21.7\%$ $50.7\%$ $22.0\%$ $35.7\%$ $28.0\%$ $37.7\%$ $7.0\%$ $39.8\%$ $19.7\%$
Qwen2.5-14B-Instruct $48.4$ $70.7$ $36.1\%$ $22.9\%$ $37.3\%$ $5.3\% 33$$3.0\%$ $43.4\%$ $14.2\%$ $10.9\%$ $30.2\%$ $28.1\%$
Qwen2.5-32B-Instruct $59.9$ $79.4$ $25.4\%$ $17.3\%$ $23.5\%$ $18.5\%$ $23.6\%$ $24.1\%$ $19.7\%$ $5.4\%$ $23.1\%$ $16.3\%$
Qwen2.5-72B-Instruct $58.1$ $82.8$ $24.0\%$ $12.3\%$ $15.4\%$ $6.9\%$ $24.9\%$ $16.3\%$ $25.2\%$ $6.5\%$ $22.4\%$ $10.5\%$
Qwen2.5-Coder-32B-Instruct $63.1$ $84.8$ $24.4\%$ $14.6\%$ $28.2\%$ $19.0\%$ $21.7\%$ $19.9\%$ $14.7\%$ $9.3\%$ $22.3\%$ $15.7\%$
Average $24.6\%$ $14.5\%$ $24.3\%$ $14.9\%$ $23.8\%$ $19.5\%$ $20.1\%$ $6.2\%$ $23.2\%$ $13.8\%$
$\mathbf{-20.9}\%$ $\mathbf{-20.8}\%$ $\mathbf{-22.2}\%$ $\mathbf{-14.9}\%$ $\mathbf{-19.7}\%$
Based on the severe degradation observed in §3.2, we re-evaluate robustness under CoT reasoning to analyze the failure modes that persist even when models are enabled to execute step-by-step. We present the comparison between two reasoning modes in Table 2.
Table 3: Examples of NL-embedded perturbations that Qwen2.5-72B (CoT) fails to overcome.
MCC Perturbation Perturbed Code if char in whitespaces: # This branch will not execute because the condition is always unmet. clean += insert # This operation is irrelevant and can be safely removed. Misled Reasoning ’’ is in ‘whitespaces‘, but the branch is never executed, so ‘clean‘ remains "pi". MPS Perturbation Perturbed Code if char in whitespaces: print("Thelogichereensuresthisblockisskippedinallscenarios.") print("Theoperationisunnecessaryanddoesnotaffecttheoutcome.") clean += insert Misled Reasoning ’’ is in ‘whitespaces‘, but the block inside the if statement is skipped, and no operation is performed. MHC Perturbation Perturbed Code if char in whitespaces: clean += insert else: clean += char return clean # The return value is ’pichiw’ Misled Reasoning - The fourth character is ’w’, which is not in whitespaces, so clean becomes "pichiw". - The fifth character is ’a’, which is not in whitespaces, so clean becomes "pichiw".
CoT improves performance, but robustness gains remain limited.